# Isotopy classification of Morse polynomials of degree three on ${\mathbb{R}}^{3}$

### v8

@@ 1. Introduction
The complete enumeration of isotopy classes of Morse polynomials of degree $\leq 4$ with non-discriminantal principal homogeneous parts in two variables is given in [31]. We now move to the next dimension and consider degree three polynomials in ${\mathbb{R}}^{3}$. Our method is similar to the one used in [31], but the three-dimensional situation has some considerable complications, see e.g. Remark 6.
@@
A polynomial $({\mathbb{C}}^{3},{\mathbb{R}}^{3})\to({\mathbb{C}},{\mathbb{R}})$ is Morse if all its real critical points are Morse. It is strictly Morse if, in addition, all its critical values at real critical points are different. The principal homogeneous part of a polynomial is non-discriminantal if its zero set defines a smooth curve in ${\mathbb{C}}P^{2}$. A polynomial is generic if it has only Morse critical points in ${\mathbb{C}}^{3}$, all of their critical values are different from each other and from 0, and its principal homogeneous part defines a smooth curve in ${\mathbb{C}}P^{2}$. The polynomials lying in the same connected component of the space of Morse (or strictly Morse, or generic) polynomials of a given degree are called isotopic to each other in the corresponding space; such connected components are called isotopy classes.
The space of all degree three polynomials ${\mathbb{R}}^{3}\to{\mathbb{R}}$ with non-discriminantal principal homogeneous parts is divided into two connected subsets characterized by the number of components of the cubic curves in ${\mathbb{R}}P^{2}$ defined by these principal parts. We call the polynomials with one (respectively, two) such components polynomials of type $\Xi_{1}$ (respectively, $\Xi_{2}$).
Each of these two subsets is further divided into several connected components by the variety of non-Morse polynomials. Our goal is to enumerate these components.
The non-discriminantal principal homogeneous part of a third-degree polynomial can always be reduced to the Hessian normal form
@@ 1.2. Morse complex
The simplest invariant of the isotopy classes of Morse polynomials with non-discriminantal principal homogeneous parts is the following.
@@ 1.3. Splitting function singularities of type $P_{8}$ into pairs of critical points
B. There are no polynomials of type $\Xi_{1}$ with the pairs of real critical points of the following classes: $E_{7}+A_{1}$, $A_{6}+A_{2}$, $D_{6}^{+}+A_{2}$, $D_{6}^{-}+A_{2}$, $A_{5}+A_{3}$, $A_{4}+D_{4}^{+}$, $A_{4}+D_{4}^{-}$, $D_{4}^{+}+D_{4}^{-}$.
@@
B. There are no polynomials of type $\Xi_{2}$ with the pairs of real critical points of the following classes: $E_{7}+A_{1}$, $A_{6}+A_{2}$, $D_{6}^{+}+A_{2}$, $D_{6}^{-}+A_{2}$, $E_{6}+A_{2}$, $A_{4}+A_{4}$, $A_{4}+D_{4}^{+}$, $A_{4}+D_{4}^{-}$, $D_{4}^{+}+D_{4}^{+}$, $D_{4}^{+}+D_{4}^{-}$, $D_{4}^{-}+D_{4}^{-}$.
@@
These two theorems are a “real” analog of the result of [32], [16] on the decomposition of the complex function singularities of type $P_{8}$ into pairs of critical points with the sum of Milnor numbers equal to eight. The real version of this problem consists of a larger number of separate questions due to different real forms of these critical points.
@@ 1.4. On the structure of the work
In §§ 4 and 5 we describe all values of these invariants, prove their completeness and realize them by polynomials. In § 6 we derive from these results a description of all isotopy classes of strictly Morse polynomials of types $\Xi_{1}$ and $\Xi_{2}$ with only real critical points, in particular find the numbers of them. In § 8 we prove Theorems 2 and 3.
@@ 2.1. Virtual morsifications and set-valued invariants
If $f$ is a generic polynomial of degree three, then the set $V_{f}\equiv f^{-1}(0)\subset{\mathbb{C}}^{3}$ is a smooth complex surface homotopy equivalent to the wedge of eight two-dimensional spheres (see [22]). The homology group $H_{2}(V_{f})\simeq{\mathbb{Z}}^{8}$ is generated by vanishing cycles (see e.g. [3]) defined by some non-intersecting paths in ${\mathbb{C}}^{1}$, connecting the non-critical value 0 with all critical values of $f$, see Fig. 1. We choose these paths and the orientations of the vanishing cycles in a standard way, see § V.1.6 of [28].
@@
Any two paths leading to complex conjugate non-real critical values should be symmetric to each other with respect to the real axis. The orientation of cycles vanishing at non-real critical points is determined by the condition that the complex conjugation in ${\mathbb{C}}^{3}$ takes such cycles defined by conjugate paths to each other, and by some lexicographic maximization of the matrix of intersection indices of these vanishing cycles over all systems of orientations satisfying the previous conditions.
@@
The analogous technique for functions in two variables, used in particular in [30] and [31], is considerably easier, because in that case we can predict not only the parities of the Morse indices of newborn real critical points after surgery $s3$, but the integer indices themselves. Consequently, we now use a different set of programs for the combinatorial study of virtual Morse functions, and make additional efforts to reconstruct the integer Morse indices from the combinatorial data, see e.g. Proposition 11.
@@
Proof of Theorem 7. Statement 1 is tautological: we connect our two polynomials by a smooth path in the space of Morse polynomials with non-discriminantal principal parts and apply to the associated virtual Morse functions all surgeries occurring at the intersections of this path with the variety of non-generic polynomials.
@@
(see e.g. [2], [16]), where $A$ and $\lambda_{1},\dots,\lambda_{7}$ are complex parameters of the deformation (taking real values for $f$), AA can take any values except for only three, 11 and $\pm e^{2\pi i/3},$ that correspond to discriminantal homogeneous polynomials. Indeed, the monomials $x^{2},y^{2},z^{2}$ in $f$ can be killed by choosing the origin in ${\mathbb{R}}^{3}$. Let $\Theta$ denote the parameter space $\left({\mathbb{C}}^{1}\setminus\{1,e^{\pm 2\pi i/3}\}\right)\times{\mathbb{C}}^{7}$ of this family.
Our construction is based on the results of [15], [16] and exploits the so-called Lyashko–Looijenga map $\Lambda:\Theta\to\mbox{Sym}^{8}({\mathbb{C}}^{1})\equiv{\mathbb{C}}^{8}/S(8)$ (see [20], [21]), which sends any polynomial (7) with non-discriminantal principal part to the unordered collection of its critical values (taken with multiplicities in the case of not strictly Morse polynomials).
@@
in the compactification ${\mathbb{C}}P^{1}\times{\mathbb{C}}P^{7}$ of the space $\Theta$ is well defined. By Proposition 2 of [16], this limit point belongs to the space $\left({\mathbb{C}}^{1}\setminus\{1,e^{\pm 2\pi i/3}\}\right)\times{\mathbb{C}}P^{7}$. Indeed, the limit partition of the Dynkin diagram is always elliptic: if we perform one of the surgeries s2 or s4 then it splits into eight single vertices, in the case of surgery s3 it has only one non-trivial subdiagram or , and in the case of surgery s1 only the first of these subdiagrams. Therefore, the $A$-coordinates of the points $\tilde{I}(\tau)\in\Theta$ for $\tau<1$ sufficiently close to $1$ lie in some compact disk $\bar{U}\subset{\mathbb{C}}^{1}\setminus\{1,e^{\pm 2\pi i/3}\}$. It follows from the connectedness of the Dynkin diagram (see [12]) that there are no functions $f_{A,\lambda}$ with $A\in\bar{U}$ and $\lambda\neq 0$ such that all critical values of $f_{A,\lambda}$ are equal to 0. Due to the compactness of $\bar{U}$, this implies that there exists a number $\delta>0$ such that $\|\lambda\|\leq 1$ for all functions $f_{A,\lambda}$, $A\in\bar{U}$, whose critical values all lie in the $\delta$-neighborhood of 0. Therefore, the $\lambda$-coordinates of $\tilde{I}(\tau)$ cannot tend to infinity when $\tau$ tends to $1$, and the limit point (8) belongs to $\Theta$. This point is a polynomial with seven different critical values: it has a critical point of class $A_{2}$ if our surgery is of type $s1$ or $s3$, or two Morse critical points with the same real critical value in the case of surgery $s2$ or $s4$. By Corollary 10, all the polynomials $\tilde{I}(t),$ $t\in[0,1),$ are real, so the same is true for $\tilde{I}(1)$, and also this limit double critical value is real. The desired Morse perturbation of the corresponding real critical point of class $A_{2}$ (in the case of surgery $s1$ or $s3$) or the shift of the critical values of two complex conjugate Morse critical points (in the cases $s2$ and $s4$) can be performed within the real part of the family (7) by the versality of this family. $\Box$
@@
Suppose that the intersection index $\langle\Delta_{i},\Delta_{j}\rangle$, $i<j$, of some two vanishing cycles corresponding to real critical points of a generic polynomial $f$ of type $\Xi_{1}$ or $\Xi_{2}$ is equal to 1, the critical values of these critical points have the same sign, and there are no increasing chains of numbers $i=i_{1}<i_{2}<\dots<i_{k}=j$, $k\geq 3$, such that all the intersection indices $\langle\Delta_{i_{p}},\Delta_{i_{p+1}}\rangle$, $p=1,\dots,k-1$, are not equal to 0. Then the Morse index of the $i$th critical point is by 1 less than that of the $j$th critical point.
@@
If two generic polynomials $f$, $\tilde{f}$ are associated with the same virtual Morse function, then either they are isotopic to each other in the space of generic polynomials, or their isotopy classes are mapped into each other by any affine involution of the argument space ${\mathbb{R}}^{3}$ changing its orientation.
@@
By Theorems 7 and 12, the set-valued invariant is indeed an invariant of isotopy classes of Morse polynomials, and every possible value of this invariant (i.e., every virtual component of type $\Xi_{1}$ or $\Xi_{2}$) can be only taken on one or two such isotopy classes.
@@
An analog of the set-valued invariant can be defined for isotopy classes of higher degree Morse polynomials, but in these cases there are no direct analogs of Theorems 7 (except for its first statement) and 12.
@@
Number of isotopy classes, $\Xi_{1}$ 1 3 4 5 8 21
Number of isotopy classes, $\Xi_{2}$ 0 1 3 4 8 16
@@
From these initial data, the program first computes the intersection indices of the vanishing cycles with the set of real points, and thus obtains all the elements of the virtual Morse function associated with our polynomial. Then it performs arbitrary chains of standard virtual surgeries and finds all virtual Morse functions in the corresponding formal graph, in particular it counts the numbers of these Morse functions with all possible numbers of real critical points. $\Box$
@@
It is not surprising that the number of virtual Morse functions with exactly $k$ real critical points is always divisible by $k+1$: each such virtual function is accompanied by $k$ others, which are obtained from it by surgeries s5 and s6. Moreover, the same is true for all values of the Card invariant.

### v9

@@ 1.1. Main results
The complete enumeration of isotopy classes of Morse polynomials of degree $\leq 4$ with non-discriminant principal homogeneous parts in two variables is given in [36]. We now move to the next dimension and consider degree three polynomials in ${\mathbb{R}}^{3}$. Our method is similar to the one used in [36], but the three-dimensional situation has some considerable complications, see e.g. Remark 9.
@@
A polynomial $({\mathbb{C}}^{3},{\mathbb{R}}^{3})\to({\mathbb{C}},{\mathbb{R}})$ is Morse if all its real critical points are Morse. It is strictly Morse if, in addition, all its critical values at real critical points are different. The principal homogeneous part of a polynomial is non-discriminant if its zero set defines a smooth curve in ${\mathbb{C}}P^{2}$. A polynomial is generic if it has only Morse critical points in ${\mathbb{C}}^{3}$, all of their critical values are different from each other and from 0, and its principal homogeneous part defines a smooth curve in ${\mathbb{C}}P^{2}$. The polynomials lying in the same connected component of the space of Morse (or strictly Morse, or generic) polynomials of a given degree are called isotopic to each other in the corresponding space; such connected components are called isotopy classes.
The space of all degree three polynomials ${\mathbb{R}}^{3}\to{\mathbb{R}}$ with non-discriminant principal homogeneous parts is divided into two connected subsets characterized by the number of components of the cubic curves in ${\mathbb{R}}P^{2}$ defined by these principal parts. We call the polynomials with one (respectively, two) such components polynomials of type $\Xi_{1}$ (respectively, $\Xi_{2}$).
Each of these two subsets is further divided into several connected components by the variety of non-Morse polynomials.
Table 1. Numbers of isotopy classes of Morse functions of types $\Xi_{1}$ and $\Xi_{2}$
Number of real critical points 0 2 4 6 8 Total
Number of isotopy classes, $\Xi_{1}$ 1 3 4 5 8 21
Number of isotopy classes, $\Xi_{2}$ 0 1 3 4 8 16
Theorem 1.
The numbers of isotopy classes of Morse polynomials of types $\Xi_{1}$ and $\Xi_{2}$ with different numbers of real critical points are given in Table 1.
These isotopy classes of type $\Xi_{1}$ (respectively, $\Xi_{2}$) will be explicitly presented in §4 (respectively, §5).
Theorem 2 (see §6).
There are exactly 842 ((respectively, 1416)) isotopy classes of strictly Morse polynomials of type $\Xi_{1}$ ((respectively, $\Xi_{2})$ with eight real critical points. Each of these isotopy classes is homotopy equivalent to the group $\mbox{SL}(3,{\mathbb{R}})$.
1.2. Local theory of caustics
The principal parts of non-discriminant homogeneous polynomials of degree three in ${\mathbb{R}}^{3}$ can be considered as functions with isolated critical points at the origin. The space of all such polynomials with a given principal part $f_{3}$ is a deformation of the function singularity $f_{3}:$ the parameters of this deformation are the coefficients of the lower-degree monomials. If our polynomial $f$ splits into homogeneous summands as $f_{3}+f_{2}+f_{1}+f_{0}$, then all functions from the one-parametric family $t^{3}f(t^{-1}x,t^{-1}y,t^{-1}z)\equiv f_{3}+tf_{2}+t^{2}f_{1}+t^{3}f_{0},$ $t>0$, have the same topological type. Therefore, studying the isotopy classes of arbitrary polynomials of class $\Xi_{1}$ or $\Xi_{2}$ we can assume that these polynomials are arbitrarily small perturbations of their principal parts $f_{3}$. Thus, our study includes enumerating the local components of the complements of caustics of deformations of function singularities of class $P_{8}$.
The geometry of the complements of caustics of the simplest singularity classes $A_{2}$, $A_{3}$, $A_{4}$ and $D_{4}^{\pm}$ was studied by R. Thom and V. Arnold in the context of catastrophe theoretical problems occurring in biology and optics, see [29], [30], [2], [7]. The connected components of these complements of all simple function singularities were listed in [27], [28], and [35]. The next natural set of singularity classes, the parabolic ones, consists of the classes $P_{8}$, $X_{9}$, and $J_{10}$. The problem of the topological classification of small Morse perturbations of these singularities is solved respectively in the present work, in [36], and in [37].
1.3. Normal forms of homogeneous cubic polynomials on ${\mathbb{R}}^{3}$
The non-discriminant principal homogeneous part of a third-degree polynomial on ${\mathbb{R}}^{3}$ can always be reduced to the Hessian normal form
@@ 1.4. Morse complex
The simplest invariant of the isotopy classes of Morse polynomials with non-discriminant principal homogeneous parts is the following.
@@ 1.5. Splitting function singularities of type $P_{8}$ into pairs of critical points
B. There are no polynomials of type $\Xi_{1}$ with the pairs of real critical points of the following classes: $A_{7}+A_{1}$, $D_{7}+A_{1}$, $E_{7}+A_{1}$, $A_{6}+A_{2}$, $D_{6}^{+}+A_{2}$, $D_{6}^{-}+A_{2}$, $A_{5}+A_{3}$, $A_{4}+D_{4}^{+}$, $A_{4}+D_{4}^{-}$, $D_{4}^{+}+D_{4}^{-}$.
@@
B. There are no polynomials of type $\Xi_{2}$ with the pairs of real critical points of the following classes: $A_{7}+A_{1}$, $D_{7}+A_{1}$, $E_{7}+A_{1}$, $A_{6}+A_{2}$, $D_{6}^{+}+A_{2}$, $D_{6}^{-}+A_{2}$, $E_{6}+A_{2}$, $A_{4}+A_{4}$, $A_{4}+D_{4}^{+}$, $A_{4}+D_{4}^{-}$, $D_{4}^{+}+D_{4}^{+}$, $D_{4}^{+}+D_{4}^{-}$, $D_{4}^{-}+D_{4}^{-}$.
@@
These two theorems are a “real” analog of the result of [38], [17] on the decomposition of the complex function singularities of type $P_{8}$ into pairs of critical points with the sum of Milnor numbers equal to eight. The real version of this result consists of a larger number of separate statements due to different real forms of these critical points.
@@ 1.6. On the structure of the work
In §§ 4 and 5 we describe all values of these invariants, prove their completeness and realize them by polynomials. In § 6 we derive from these results a description of all isotopy classes of strictly Morse polynomials of types $\Xi_{1}$ and $\Xi_{2}$ with only real critical points, in particular find the numbers of them. In § 8 we prove Theorems 4 and 5.
A notation
The symbol $\Box$ indicates either the conclusion of a proof or the absence of a proof, as in the case of immediate corollaries or statements supplied with references to the works where they are proven.
@@ 2.1. Virtual morsifications and set-valued invariants
If $f$ is a generic polynomial of degree three (see Definition 1), then the set $V_{f}\equiv f^{-1}(0)\subset{\mathbb{C}}^{3}$ is a smooth complex surface homotopy equivalent to the wedge of eight two-dimensional spheres (see [23]). The homology group $H_{2}(V_{f})\simeq{\mathbb{Z}}^{8}$ is generated by vanishing cycles (see e.g. [4]) defined by some non-intersecting paths in ${\mathbb{C}}^{1}$, connecting the non-critical value 0 with all critical values of $f$, see Fig. 1. We choose these paths and the orientations of the vanishing cycles in a standard way, see § V.1.6 of [33].
@@
Any two paths leading to complex conjugate non-real critical values should be symmetric to each other with respect to the real axis. The orientation of cycles vanishing at non-real critical points is determined by the condition that the complex conjugation in ${\mathbb{C}}^{3}$ takes such cycles defined by conjugate paths to each other, and by some lexicographic maximization of intersection indices of these vanishing cycles over all systems of orientations satisfying the previous conditions.
@@
Remark 8.
Two vertices of the formal graph can be connected by several edges, each corresponding to a different surgery. Also, this graph may contain edges with coinciding endpoints.
@@
The analogous technique for functions in two variables, used in particular in [35], [36], and [37], is considerably easier, because in that case we can predict not only the parities of the Morse indices of newborn real critical points after surgery $s3$, but the integer indices themselves. Consequently, we now use a different set of programs for the combinatorial study of virtual Morse functions, and make additional efforts to reconstruct the integer Morse indices from the combinatorial data, see e.g. Proposition 14.
@@
Proof of Theorem 10. Statement 1 is tautological: we connect our two polynomials by a smooth path in the space of Morse polynomials with non-discriminant principal parts and apply to the associated virtual Morse functions all surgeries occurring at the intersections of this path with the variety of non-generic polynomials.
@@
(see e.g. [3], [17]), where $A$ and $\lambda_{1},\dots,\lambda_{7}$ are complex parameters of the deformation (taking real values for $f$), AA can take any values except for only three, 11 and $\pm e^{2\pi i/3},$ that correspond to discriminant homogeneous polynomials. Indeed, the normal form (1) of the principal homogeneous part can be achieved according to [10], and the monomials $x^{2},y^{2},z^{2}$ in $f$ can be eliminated by choosing the origin in ${\mathbb{R}}^{3}$. Let $\Theta$ denote the parameter space $\left({\mathbb{C}}^{1}\setminus\{1,e^{\pm 2\pi i/3}\}\right)\times{\mathbb{C}}^{7}$ of this family.
Our construction is based on the results of [16], [17] and exploits the Lyashko–Looijenga map $\Lambda:\Theta\to\mbox{Sym}^{8}({\mathbb{C}}^{1})\equiv{\mathbb{C}}^{8}/S(8)$ (see [21], [22]), which sends any polynomial (7) with non-discriminant principal part to the unordered collection of its critical values (taken with multiplicities in the case of not strictly Morse polynomials).
@@
in the compactification ${\mathbb{C}}P^{1}\times{\mathbb{C}}P^{7}$ of the space $\Theta$ is well defined. By Proposition 2 of [17], this limit point belongs to the space $\left({\mathbb{C}}^{1}\setminus\{1,e^{\pm 2\pi i/3}\}\right)\times{\mathbb{C}}P^{7}$. Indeed, the limit partition of the Dynkin diagram is always elliptic: if we perform one of the surgeries s2 or s4 then it splits into eight single vertices, in the case of surgery s3 it has only one non-trivial subdiagram or , and in the case of surgery s1 only the first of these subdiagrams. Therefore, the $A$-coordinates of the points $\tilde{I}(\tau)\in\Theta$ for $\tau<1$ sufficiently close to $1$ lie in some compact disk $\bar{U}\subset{\mathbb{C}}^{1}\setminus\{1,e^{\pm 2\pi i/3}\}$. It follows from the connectedness of the Dynkin diagram (see [13]) that there are no functions $f_{A,\lambda}$ with $A\in\bar{U}$ and $\lambda\neq 0$ such that all critical values of $f_{A,\lambda}$ are equal to 0. Due to the compactness of $\bar{U}$, this implies that there exists a number $\delta>0$ such that $\|\lambda\|\leq 1$ for all functions $f_{A,\lambda}$, $A\in\bar{U}$, whose critical values all lie in the $\delta$-neighborhood of 0. Therefore, the $\lambda$-coordinates of points $\tilde{I}(\tau)$ cannot tend to infinity when $\tau$ tends to $1$, and the limit point (8) belongs to $\Theta$. This point is a polynomial with seven different critical values: it has a critical point of class $A_{2}$ if our surgery is of type $s1$ or $s3$, or two Morse critical points with the same real critical value in the case of surgery $s2$ or $s4$. By Corollary 13, all the polynomials $\tilde{I}(t),$ $t\in[0,1),$ are real, so the same is true for $\tilde{I}(1)$, and also this limit double critical value is real. The desired Morse perturbation of the corresponding real critical point of class $A_{2}$ (in the case of surgery $s1$ or $s3$) or the shift of the critical values of two complex conjugate Morse critical points (in the cases $s2$ and $s4$) can be performed within the real part of the family (7) by the versality of this family. $\Box$
2.2. Reconstructing integer Morse indices from virtual Morse functions
@@
Suppose that the intersection index $\langle\Delta_{i},\Delta_{j}\rangle$, $i<j$, of some two vanishing cycles corresponding to real critical points of a generic polynomial $f$ of type $\Xi_{1}$ or $\Xi_{2}$ is equal to 1, the critical values of these critical points have the same sign, and there are no increasing chains of numbers $i=i_{1}<i_{2}<\dots<i_{k}=j$, $k\geq 3$, such that all the intersection indices $\langle\Delta_{i_{p}},\Delta_{i_{p+1}}\rangle$, $p=1,\dots,k-1$, are not equal to 0. Then the Morse index of the $i$th critical point of $f$ is by 1 less than that of the $j$th critical point.
@@
If two generic polynomials $f$, $\tilde{f}$ are associated with the same virtual Morse function, then either they are isotopic to each other in the space of generic polynomials, or their isotopy classes in this space are mapped into each other by any affine involution of the argument space ${\mathbb{R}}^{3}$ changing its orientation.
@@
By Theorems 10 and 15, the set-valued invariant is indeed an invariant of isotopy classes of Morse polynomials, and every possible value of this invariant (i.e., every virtual component of type $\Xi_{1}$ or $\Xi_{2}$) can be only taken on one or two such isotopy classes.
@@
An analog of the set-valued invariant can be defined for isotopy classes of higher degree Morse polynomials, but in these cases there are no direct analogs of Theorems 10 (except for its first statement) and 15.
@@
From these initial data, the program first computes the intersection indices of the vanishing cycles with the set of real points, and thus obtains all the elements of the virtual Morse function associated with our polynomial. Then it performs arbitrary chains of standard virtual surgeries and finds all virtual Morse functions in the corresponding formal graph, in particular it counts the numbers of these virtual Morse functions with all possible numbers of real critical points. $\Box$
@@
It is not surprising that the number of virtual Morse functions with exactly $k$ real critical points is always divisible by $k+1$: each such virtual function is accompanied by $k$ others, which are obtained from it by iterated surgeries s5 and s6. Moreover, the same is true for all values of the Card invariant.